3.   Given three subroutines of 550, 290, and 600 words each, if segmentation is used then the total memory needed is the sum of the three sizes (if all three routines are loaded). However, if paging is used then some storage space is lost because subroutines rarely fill the last page completely, and that results in internal fragmentation. Determine the total amount of wasted memory due to internal fragmentation when the three subroutines are loaded into memory using each of the following page sizes:

a.   100 words

b.   600 words

c.   700 words

d.   900 words

hint: you may want to consider making a table with job size and page size on the x and y axis to answer the questions.
Answer #1

job size 550 +290+ 600 = 1440 words

a. page size =100 words   

internal fragmentation =(100*15 )-1440= 60 words

b. page size = 600 words

internal fragmentation = (600*3)- 1440 = 360 words

c. page size = 700 words

internal fragmentation = (700*3) -1440 = 660 words

d. page size = 900 words

internal fragmentation = (900*2) -1440 = 360 words